Previous

Calendar

January

2008

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

30

JD364

31

JD365

1

JD1
CU Holiday (New Years Day)

2

JD2
Site Visit (JH, RJ): Details

at 4:02 MST, russter2 locks up and the data system goes down.

at 10:40 MST, russter2 rebooted and data system back up.

at 13:20 MST, russter2 locks up (again) and the data system goes down.

3

JD3 (warm)
Site Visit (SpB): Details , Photos (17)

at 10:15 MST, russter2 rebooted (data system not started).

at 17:00 MST, russter2 removed from trailer and brought to campus.

Prius Rental

N2 @ 1150 psi

span @ 200 psi

downloaded from all loggers except Noah's logger

4

JD4 (windy)
Site Visit (SpB, NT, SL): Details

at 16:30 MST, russter2 returned to trailer, rebooted, and data system re-started.

Jeep Rental

5

JD5

6

JD6
JB returns

7

JD7

8

JD8

9

JD9 (cold, breezy, snowing)
Site Visit (JB, SL): Details

re-started the west robot sampling every other week

swapped pcmcia cards at the mrs logger

@ 1:32 MST, stopped EC system GAST pump, swapped 2-micron Nupro filters at the EC inlet and re-started the pump @ 1:53 MST

swapped desiccant at the LI-6251 between 2:04 and 2:24 MST

retired Scott-Marrin span cylinder CC66823 @ 150 psi (hauled back to trailer), started Airgas CC18996 @ 400 psi

@ ~3:45 MST saw a female moose and her calf along the main trail between C1 and the tower

10

JD10

11

JD11

12

JD12

13

JD13

14

JD14 (calm and warm early, wind picked up in the afternoon)
Site Visit (JB, SpB): Details , Photos (17)

CU classes begin for spring semester

downloaded from all loggers except UC and UCB (and didn't swap pcmcia cards at the mrs logger--do this next visit)

removed the 21.5m CSAT3 sonic anemometer (sn 0226) to return to Campbell for its first re-cal in ten years! swapped it with a loaner CSAT3 sonic from EOL (sn 0536)

15

JD15

16

JD16

17

JD17
Campbell received and started the re-cal process on our 21.5m csat3 this morning

18

JD18
SpB Off (sick day)

19

JD19

20

JD20

21

JD21
CU holiday (Martin Luther King, Jr. Day)

22

JD22

23

JD23
JB has 10 AM dentist appointment

24

JD24

25

JD25
Site Visit (JB, NT): Details

re-calibrated csat3 returned to CU

hauled new span cylinder (CC199497) to the tower

hauled old N2 cylinder (CC199706, 100 psi) to trailer

current N2 @ 400 psi

current span @ 250 psi

swapped 4 flasks as west robot, changed logger program to account for correct fudge day (now=5, previously=6)

downloaded from soil, tc, uc, ucb, and swapped pcmcia cards at mrs logger; all data files given correct post-processing name and moved to /data/projects/NIWOT/data_logger/

dug out the lowest profiler inlets at the mini tower and dug out the uc/ucb logger boxes

see details for snow depths

26

JD26

27

JD27

28

JD28

29

JD29 (-20 C, peak gusts 25 m/s)
Site Visit (JB, SpB): Details , Photos (11)

re-deployed the re-calibrated CSAT3 sonic (sn 0226) at 2.5 m on the mini tower, removed the 2.5 m CSAT sonic (sn 0438) to send back for re-cal

JB and SpB helped Duane/Dwayne? load NT's snow blower onto the snow cat

SpB flagged off an area containing subterranean sensors around the mini tower...fortify this with rope/string next visit

30

JD30

31

JD31

1

JD32
Site Visit (JB, SLS, NT): Details

switched the SDM setting on the re-calibrated sonic from 3 to 4; this got it working fine

reinforced the flagging around the mini tower

power cylcled the gmp343 probe and got it back on scale in cockpit

span @ 200 psi; N2 @ 190 psi; retired the N2 (CC198536 @ 190 psi, started N2 CC125555 @ 400 psi -- this was Jia's summer pressure bomb cylinder)

see details for snow depths

2

JD33